OpenStudy (qqstory):

How does sec^2(x) / sec^4(x) = cos^2(x)?

OpenStudy (qqstory):

sec^-2(x) = 1 / 1 / cos^2(x)

OpenStudy (qqstory):

\[\frac{ 1 }{ 1 } * \frac{ \cos^2 x }{ 1 }\]

Directrix (directrix):